### v2

# WaveDiffusion: Exploring Full Waveform Inversion via Joint Diffusion in the Latent Space

Abstract
Full Waveform Inversion (FWI) reconstructs high-resolution subsurface velocity maps from seismic waveform data governed by partial differential equations (PDEs). Traditional machine learning approaches frame FWI as an image-to-image translation task, mapping seismic data to velocity maps via encoder-decoder architectures. In this paper, we revisit FWI from a new perspective: generating both modalities simultaneously. We found that both modalities can be jointly generated from a shared latent space using a diffusion process. Remarkably, our jointly generated seismic-velocity pairs inherently satisfy the governing PDE without requiring additional constraints. This reveals an interesting insight: the diffusion process inherently learns a scoring mechanism in the latent space, quantifying the deviation from the governing PDE. Specifically, the generated seismic-velocity pairs with higher scores are closer to the solutions of the governing PDEs. Our experiments on the OpenFWI dataset demonstrate that the generated seismic-velocity pairs not only yield high fidelity, diversity and physical consistency, but also can serve as effective augmentation for training data-driven FWI models.
Full Waveform Inversion, Joint Diffusion, Denoising, Deviation from PDE
1 Introduction
Subsurface imaging is critical in scientific and industrial applications, including earthquake monitoring (Virieux et al., 2017; Tromp, 2020), greenhouse gas storage (Li et al., 2021; Wang et al., 2023b), medical imaging (Guasch et al., 2020; Lozenski et al., 2024), and oil and gas exploration (Virieux & Operto, 2009; Wang & Alkhalifah, 2018). At its core, subsurface imaging involves reconstructing velocity maps that describe the propagation speed of seismic waves, governed by the acoustic wave equation:
$\frac{\partial^{2}s(\mathbf{x},t)}{\partial t^{2}}=v^{2}(\mathbf{x})\nabla^{2}s(\mathbf{x},t)+f(\mathbf{x_{s}},t),$ (1)
where $s(\mathbf{x},t)$ is the seismic data, $v(\mathbf{x})$ is the velocity model, $\nabla^{2}$ is the Laplacian, and $f(\mathbf{x_{s}},t)$ is the source term. Full Waveform Inversion formulates this task as an inverse problem: given observed seismic data $s(\mathbf{x},t)$, the goal is to estimate the underlying velocity map $v(\mathbf{x})$.
Recently, machine learning-based approaches (Wu & Lin, 2019; Zhang et al., 2019; Sun & Demanet, 2020; Feng et al., 2021) have been introduced to address the limitations (computational cost, cycle skipping, etc) of traditional FWI methods (Plessix, 2006). These machine learning-based methods treat FWI as an image-to-image translation task and use neural networks to directly map from seismic data to velocity maps. In particular, they encode the seismic data $s$ into a latent space $z=E_{\theta}(s)$ from which the velocity map is decoded by $v=D_{\omega}(z)$, where $\theta$ and $\omega$ are the learnable parameters in the encoder and decoder, respectively.
This paper is motivated by an investigation of the latent space $z$. Our preliminary study reveals that randomly sampling from $z$ and computing the corresponding velocity map $v=D(z)$ and seismic reconstruction $\hat{s}=E^{-1}(z)$ (via an additional learned decoder) often results in pairs that do not satisfy the governing PDE in Equation 1. This finding suggests that only a sparse subspace of $z$ corresponds to valid PDE solutions. Naturally, this raises the question: Can we systematically distinguish latent space representations that adhere to the governing PDE from those that do not?
Surprisingly, we found that this can be achieved by applying diffusion models. After training a diffusion network in the shared latent space, we observe that the diffusion process transforms arbitrary latent points into those that correspond to physically valid solutions. This reveals an intriguing insight: the diffusion model automatically learns to score the latent space based on its deviation from the governing PDE. Specifically, higher scores are assigned to latent points from which valid seismic-velocity pairs can be decoded.
Without additional constraints, our framework results in a latent diffusion model that can jointly generate seismic-velocity pairs that approximately satisfy the governing PDE. Experiments on the OpenFWI dataset demonstrate that the generated samples exhibit high fidelity and diversity, providing a valuable source of augmented data for training conventional image-to-image translation models like InversionNet.
The key contribution of this paper is introducing a new generative perspective for FWI: the deviation from the solutions of the governing PDE can be modeled by a diffusion process in the latent space. We believe this insight deepens our understanding of deep learning-based FWI models and bridges generative AI with physics-based modeling.
2 WaveDiffusion: Our Method
In this section, we present WaveDiffusion, a framework designed to explore and refine the latent space extracted by encoder-decoder models for FWI. Specifically, we aim to address two fundamental questions: (1) Do all the points in the latent space constructed by traditional image-to-image mapping models satisfy the governing PDE? (2) If not, can we distinguish latent samples $\{z\}$ that satisfy the governing PDE from those that do not? To investigate these questions, we formulate a two-stage framework: first, we construct the latent space using an encoder-decoder model, then we refine it through a diffusion process to ensure adherence to the governing PDE.
2.1 Stage 1: Encoder-Decoder and Reconstruction
Extending encoder-decoder by adding reconstruction branch: To enhance the interpretability of the latent space, we extend conventional encoder-decoder FWI models by incorporating an additional seismic decoder branch. This modification enables the simultaneous reconstruction of both seismic and velocity maps from a shared latent representation, facilitating a deeper analysis of its structure. This enables the analysis of whether all latent space points satisfy the PDE. To facilitate the second-stage diffusion process, we incorporate vector quantizations in the latent space.
Achieving comparable performance in FWI: We evaluate the inversion performance of our vector-quantized encoder-decoder model against BigFWI-B (Jin et al., 2024), a state-of-the-art InversionNet trained on the full OpenFWI dataset. This provides a fair comparison, as the training sample volume and model parameter size of BigFWI-B (24M) approximately align with our Stage 1 encoder-decoder model (19M).
Figure 1 shows that our encoder-decoder model achieves competitive performance relative to BigFWI-B across all datasets, outperforming BigFWI-B in six out of ten datasets (marked with yellow stars).
Figure 1: Comparison of inversion performance. Mean Absolute Error (MAE) across various OpenFWI datasets for encoder-decoder-based models and diffusion refinements. Our models perform competitively against BigFWI-B, with diffusion providing slight refinements. Yellow stars indicate datasets where our model outperforms the BigFWI-B baseline.
Definition of deviation from PDE: To quantitatively investigate whether randomly sampled $z$ points satisfy the PDE, we measure the deviation of the generated seismic-velocity pairs from the governing PDE. Specifically, for a randomly sampled latent representation $z$, we decode the seismic data $\hat{s}$ and velocity map $\hat{v}$. Using a finite difference solver, we compute the ground truth seismic $s_{\hat{v}}$ for the generated $\hat{v}$. The deviation from the PDE is quantified as the L2 distance $\|\hat{s}-s_{\hat{v}}\|_{2}$.
Most latent points do not correspond to PDE solutions: The deviation of decoded modalities from randomly sampled $z$ points that are usually far away from trained latent points, reaches an average L2 distance of $0.013$ (Figure 4 backward step 0), which is eight times higher than the deviation ($0.0016$, Figure 4 forward step 0) of reconstructed modalities from the training set (trained latent points). Figure 2 visually compares the generated samples from Stage 1 (row 1) and the original OpenFWI dataset (row 3). The encoder-decoder provides coarse approximations of PDE solutions but does not fully satisfy the wave equation.
Figure 2: Comparison between generated and original samples: Examples of generated pairs by the (row 1) encoder-decoder and (row 2) joint diffusion. Row 3 shows an original OpenFWI example. The encoder-decoder-generated samples lack the physical relationships governed by the wave equation, while diffusion refines them to better satisfy the PDE.
Thus, encoding and decoding through this latent space alone does not inherently ensure physical validity. Our experiments show that most generated samples fail to satisfy the governing PDE, meaning only a sparse subspace of $z$ corresponds to valid PDE solutions.
2.2 Stage 2: Joint Diffusion in the Latent Space
Refining the latent space with diffusion: Since random sampling in the latent space does not guarantee adherence to the PDE, we use a joint diffusion model to progressively transform arbitrary latent points into valid ones. This process follows a standard forward-backward diffusion formulation:
Forward process: Gaussian noise is added to the latent vector $z$, creating a noisy representation: $z_{t}=z_{t-1}+\epsilon_{t},\quad t=1,\dots,T$, where $\epsilon_{t}$ is the noise applied at step $t$, with in total $T$ steps.
Backward process: The noisy latent vector is progressively denoised through the reverse diffusion process: $z_{t-1}=z_{t}-\gamma_{t}\nabla_{z_{t}}\mathcal{L}(z_{t},t)$, where $\gamma_{t}$ is a step size, and $\mathcal{L}$ is the denoising objective defined as:
$\mathcal{L}=\mathbb{E}_{z_{0},\epsilon\sim\mathcal{N}(0,I),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t)\|^{2}_{2}\right].$ (2)
This follows the standard formulation of denoising score matching for latent diffusion models (Rombach et al., 2022). Here, $\epsilon_{\theta}(z_{t},t)$ is the predicted noise estimated by the learned model. The loss function minimizes the discrepancy between the true noise $\epsilon$ and the model’s predicted noise $\epsilon_{\theta}(z_{t},t)$, guiding the model to iteratively refine the noisy latent variable back to a valid solution in the latent space.
Once trained, the model can generate new seismic-velocity pairs that satisfy the wave equation by sampling latent vectors from a Gaussian distribution and refining them via the learned backward process: (1) Sample a latent vector $z_{t}$ from a standard Gaussian distribution: $z_{t}\sim\mathcal{N}(0,I)$. (2) Pass $z_{t}$ through the backward denoising steps: $z_{t-1}=\mathcal{L}(z_{t}),\quad t=T,\dots,1$. (3) Decode $z_{0}$ back into seismic data and velocity maps: $\hat{s}=D_{\text{s}}(z_{0}),~{}\hat{v}=D_{\text{v}}(z_{0})$.
2.3 Inspecting Diffusion Process
To analyze the role of diffusion, we measure the deviation of generated seismic-velocity pairs from the governing PDE at each diffusion step using the L2 distance between $\hat{s}$ and $s_{\hat{v}}$. Figure 3 visualizes this process, while Figure 4 presents the statistical evaluation of 13,200 generated pairs at each diffusion step.
Figure 3: Visualization of deviation from PDE during diffusion. Seismic data comparison of a CVB example at different timesteps during the forward (left half) and backward diffusion processes (right half). Rows 1-3 show generated seismic channels $\hat{s}$ by the joint diffusion model. Row 4 shows the generated velocity map $\hat{v}$. Rows 5-7 show the ground truth seismic data $s_{\hat{v}}$ calculated for the generated $\hat{v}$. Row 8 shows the deviation from the PDE, visualized as the channel-stacked difference between $s_{\hat{v}}$ and $\hat{s}$. Noise increases the deviation during the forward diffusion, and the reverse process reduces discrepancies.
Deviation increases by adding noise: In the left half of Figure 3, during the forward diffusion process, as noise is added, the seismic data $\hat{s}$ generated by the joint diffusion model diverges more from the ground truth $s_{\hat{v}}$. This divergence, shown as the channel-stacked difference in the last row, reflects the increasing deviation from PDE as the noise level rises. Meanwhile, the decoded modalities show distortion in structures. The statistical evaluation in Figure 4 illustrates how the deviation increases with noise in the forward process, from the initial deviation 0.0016 (trained latent points) to 0.0121 (pure noise points).
Deviation decreases by denoising: In contrast, the right half of Figure 3 shows the backward process, where noise is progressively removed, reducing the deviation and refining the seismic-velocity pairs toward solutions with smaller stacked seismic difference and restored structures. Similarly the statistical evaluation in Figure 4, the deviation decreases as the generated pairs are refined into physically valid solutions from an averaged L2 distance of 0.013 back to 0.002, confirming the model’s ability to refine latent points toward physically consistent solutions.
Figure 4: Deviation from the governing PDE. The L2 distance is calculated between (a) generated seismic data $\hat{s}$ and (b) ground truth seismic data $s_{\hat{v}}$ calculated for the generated velocity map $\hat{v}$ using a finite difference solver.
Our results reveal that only a small subspace of latent points satisfies the PDE, while diffusion refines high-deviation points into physically valid solutions. This suggests that diffusion implicitly scores the latent space, moving it toward PDE-compliant solutions. The diffusion process effectively transforms the deterministic PDE problem into a stochastic differential equation (SDE), allowing exploration of the solution space. This insight provides a new perspective to bridge AI-driven generative modeling with physical principles.
3 Experiments
In this section, we present the experimental evaluation of the proposed WaveDiffusion framework. We conduct experiments on the full OpenFWI dataset to evaluate the model’s performance in generating physically consistent seismic data and velocity maps. We assess the model’s FID scores and show its ability to improve FWI results. Then, we compare the results of training the state-of-the-art models such as BigFWI using the jointly generated dataset against the original benchmark (Jin et al., 2024). In the Appendix, we further introduce an experiment to demonstrate how the joint diffusion model compares to separately trained diffusion models in generating seismic and velocity modalities.
3.1 Dataset and Training Setup
In the experiments, we evaluate two different configurations for the latent diffusion models. Specifically, we test (1) two separate VQ codebooks for the two modalities and (2) a single shared VQ codebook for both modalities.
We evaluate the performance of our WaveDiffusion framework using the OpenFWI dataset, a comprehensive benchmark collection comprising 10 subsets of realistic synthetic seismic data paired with subsurface velocity maps, specifically designed for FWI tasks. These subsets represent diverse geological structures, including curved velocity layers, flat velocity layers, and flat layers intersected by faults, among others, allowing for an extensive evaluation of our model’s robustness and generalization capability.
Our experiments utilized all 10 subsets (Fault, Vel, and Style Families) with over 400K training data pairs to ensure a thorough assessment of WaveDiffusion. We performed training on the combined dataset comprising all subsets to assess the model’s ability to generalize across a variety of geological configurations and scenarios. We trained our models on 128 NVIDIA GH200 GPUs. Network details and training hyperparameters used are provided in Appendix A.1.
3.2 Evaluating Generated Samples with FID
The Stage 1 model of our WaveDiffusion framework produces coarse approximations of seismic-velocity pairs. The Stage 1 model without a diffusion process yielded an FID score of 14,207.14 for the randomly sampled $z$ vectors and their corresponding decoded velocity maps and 871.31 for the decoded seismic data using an Inception-v3 model pre-trained on ImageNet (Szegedy et al., 2016). A visualization example is shown in Figure 2 row 1. These high FID scores suggest that the encoder-decoder architecture, while generating plausible shapes, does not adhere closely to the true data distribution. The large disparity between seismic and velocity FID scores indicates that the generated modalities deviate more from the physical relationships governed by the wave equation as coarse approximations of the PDE solutions.
Table 1: FID scores. The FID scores of the two modalities velocity maps $v$ and seismic data $s$ for two model settings.
[…]
Velocity FID 665.75 260.33
Seismic FID 20.94 5.67
The joint diffusion model in Stage 2 is used to refine the coarse generations into physically consistent seismic-velocity pairs. We also evaluate the FID scores of both modalities generated by the joint diffusion models with different vector quantization strategies. We generated 375,000 samples for each configuration, which matched the scale of the OpenFWI training set. As shown in Table 1, among the tested configurations, the joint diffusion model with a single shared VQ layer achieved the lowest FID scores, recording 260.33 for velocity and 5.67 for seismic data, indicating a stronger latent space connection between the two modalities.
Figure 5: t-SNE visualization of features for both real (blue) and generated (orange) data. Results of seismic data (col 1) and velocity map (col 2) for two joint diffusion models.
Figure 5 presents a t-SNE visualization of the feature representations extracted from Inception V3 for both real and generated data, which are used to compute the FID score. The visualization provides insight into the distributional alignment between real and synthesized samples in the feature space. A closer overlap between the two distributions indicates better fidelity of the generated data. This figure serves as a qualitative complement to the FID scores, illustrating how well the joint diffusion model captures the underlying data distribution.
Figure 6 visualizes results from the model with two VQs, while samples from the model with one VQ are shown in Appendix A.2. In the top row of each figure, we select samples structurally similar to FlatVel-B (FVB), where seismic inputs exhibit perfect symmetry, and the generated velocity maps maintain this symmetry, demonstrating the model’s ability to respect geometric constraints. Beyond reproducing individual OpenFWI subsets, we also observe cases where features from multiple datasets are fused, as seen in the ninth row of Figure 6, showcasing the diversity and fidelity of the generated seismic-velocity pairs. These results confirm that the joint diffusion model generalizes well across different data distributions, effectively capturing structural coherence and producing reliable outputs for seismic data-velocity map generation.
Figure 6: Generated examples from the joint diffusion model with two VQs.
3.3 Diffusion Improves FWI Performance
This experiment evaluates the inversion performance of our models of the two stages, specifically the encoder-decoder with two individual VQ layers (Stage 1) and its corresponding latent diffusion refinements (Stage 2). We compare these models against the BigFWI-B model, which serves as a fair baseline since its training sample volume and model parameter size approximately align with our encoder-decoder model.
For this experiment, the Stage 1 model functions as an image-to-image translation network, analogous to BigFWI-B. In contrast, the diffusion model in Stage 2 refines the latent representation using the last 10 backward denoising steps, aiming to improve reconstruction accuracy. The refinement process is performed as follows: at each denoising step, we randomly sample 100 new latent vectors $z_{t}$ and decode them into seismic data $\hat{s}$. Each decoded $\hat{s}$ is compared with the corresponding input seismic data $s$ by computing the L2 distance. The sample with the lowest L2 distance is selected as the current $z_{t}$ for the next backward denoising step. This iterative refinement continues until reaching $z_{0}$, at which point the final reconstructed seismic-velocity pair is obtained.
Improvement by diffusion: Our results, as shown in Figure 1 for metrics comparison, demonstrate that the Stage 1 model achieves inversion performance comparable to BigFWI-B across all datasets. Additionally, applying the diffusion model to refine the latent space leads to slight but consistent improvements in reconstruction accuracy. Visualization of examples can be found in Figure 7. These findings suggest that the latent representations learned by the Stage 1 model already encode meaningful physical structures, and the diffusion process further enhances them by guiding reconstructions toward physically valid solutions.
Figure 7: Inversion performance visualization. Inverted examples from CVB (row 1), FFB (row 2), and FVB (row 3) subsets. (Col 1) ground truth; (col 2) 2VQ without diffusion; (col 3) 2VQ with diffusion; (col 4) BigFWI-B.
For a more in-depth analysis of numerical comparisons, including RMSE, MAE, and SSIM metrics, we refer readers to Appendix A.3.
3.4 Training Data Augmentation
Table 2: Quantitative results of InversionNet performance (MAE) training on pure generated data. Performance of InversionNet trained on jointly generated and original OpenFWI data.
Dataset FVA FVB CVA CVB FFA FFB CFA CFB SA SB
2VQ-Generated Data 0.0560 0.1134 0.1001 0.2130 0.0636 0.1453 0.0706 0.2065 0.1103 0.1061
1VQ-Generated Data 0.0579 0.1257 0.0952 0.2030 0.0632 0.1395 0.0730 0.1997 0.1103 0.1064
BigFWI-B 0.0055 0.0233 0.0343 0.0933 0.0106 0.0710 0.0167 0.1245 0.0514 0.0553
Table 3: Quantitative results of InversionNet performance (MAE) for partial real data and partial real + generated data. Performance of InversionNet trained on partial OpenFWI data and partial real data plus generated data.
Dataset FVA FVB CVA CVB FFA FFB CFA CFB SA SB
Gen + 10%Real 0.0192 0.0687 0.0675 0.1652 0.0262 0.1140 0.0400 0.1817 0.0862 0.0808
Gen + 1%Real 0.0386 0.1107 0.0846 0.1987 0.0462 0.1320 0.0588 0.1957 0.0968 0.0908
10%Real 0.0328 0.0978 0.0934 0.2294 0.0453 0.1496 0.0616 0.2116 0.1121 0.0984
1%Real 0.0983 0.2445 0.1507 0.3402 0.1140 0.1997 0.1339 0.2550 0.1670 0.1317
We evaluate the effectiveness of WaveDiffusion-generated samples in augmenting training data, assessing how well they supplement the original OpenFWI dataset.
Real data vs. generated data: To examine the standalone quality of generated data, we train InversionNet (Wu & Lin, 2019) exclusively on WaveDiffusion-generated samples, ensuring the dataset matches the scale of the OpenFWI training set. The model is then tested on the OpenFWI test set, and results are compared against the state-of-the-art BigFWI-B (Jin et al., 2024) baseline. As shown in Table 2, models trained purely on generated data underperform compared to those trained on real data, indicating a gap in the fidelity of the generated samples.
Partial real data vs. partial real + gen: We further investigate the impact of mixing generated data and partial real data for training, focusing on samples generated by the model with one shared VQ. Specifically, we evaluate two settings: Gen+$n\%$Real, where generated data is combined with a small portion ($10\%$ or $1\%$) of real data, and $n\%$Real, where models are trained exclusively on the same small subset of real data.
Table 3 presents the results for both settings. Incorporating even a small fraction of real data (Gen+$n\%$Real) significantly improves performance. Gen+10%Real demonstrates a substantial improvement over using 10%Real alone, with metrics approaching the BifFWI-B. Similarly, Gen+1%Real outperforms 10%Real on most datasets, particularly on more complex data. These findings highlight the effectiveness of augmenting generated data with even a small amount of real data. In contrast, the $n\%$Real models exhibit the weakest performance, emphasizing the limitations of relying solely on a small dataset for effective training.
Figure 8: InversionNet performance visualization. Examples from CVB subset.
A prediction visualization is shown in Figure 8. These results demonstrate that while generated samples can effectively supplement small datasets, the inclusion of even a small portion of real data is crucial for achieving optimal results. More prediction visualizations are shown in Appendix A.4.
4 Related works
Traditional physics-based FWI: Traditional FWI methods aim to reconstruct subsurface velocity models by iteratively minimizing the difference between observed and simulated seismic data, typically using gradient-based optimization methods. The key challenge lies in solving the wave equation, which governs wave propagation through the Earth. While effective, these methods are computationally expensive and sensitive to factors such as the quality of the initial velocity model, noise in the data, and cycle-skipping issues—where the inversion algorithm converges to incorrect solutions due to poor starting models or insufficient low-frequency data (Tarantola, 1984; Virieux & Operto, 2009). Techniques such as adaptive waveform inversion (Warner & Guasch, 2016) and multiscale FWI (Bunks et al., 1995) have been developed to reduce the risk of cycle-skipping and improve convergence by progressively introducing higher-frequency data. These techniques frame FWI as a conditional generation problem, relying on physical equations as the computational foundation (Virieux et al., 2017; Warner & Guasch, 2016).
Data-driven approaches to FWI: In recent years, machine learning approaches have been increasingly explored for FWI. Convolutional Neural Networks (CNNs) have shown promise in learning image-to-image mappings from seismic data to velocity models, bypassing the need for iterative solvers. Encoder-decoder architectures, such as those used in InversionNet (Wu & Lin, 2019) and VelocityGAN (Zhang et al., 2019), have demonstrated the ability to predict velocity maps from seismic data while reducing computational costs by learning implicit relationships between the two modalities. Richardson’s work (Richardson, 2018) further illustrated that deep learning models could predict velocity models efficiently. However, these approaches are still treating the FWI problem as an image-to-image translation task or a conditional generation problem (Zhu et al., 2019; Wang et al., 2023a). Recent work on neural operators (Li et al., 2020, 2023) offers a more flexible approach by learning operators that map between the two modalities in a revered direction, i.e., predicting seismic data given velocity maps. While neural operator methods demonstrate a strong capacity in modality mapping, their reliance on direct image-to-image translation prevents them from capturing the intrinsic physical constraints embedded in the latent space.
Generative models in FWI: Generative models, particularly Generative Adversarial Networks (GANs) and their variants, have emerged as alternatives to traditional CNN-based methods for FWI. These models aim to learn the latent representations of seismic data and velocity models, enabling the generation of synthetic training data or even direct inversion (Goodfellow et al., 2020). Vector Quantized GANs (VQGANs) (Esser et al., 2021), in particular, have been explored for their ability to generate high-quality modalities, such as images, audios, videos, etc. Such models can be tuned for imaging one physical modality (e.g., velocity) given another (e.g., seismic) (Zhang et al., 2019).
Recent work has focused on Latent Diffusion Models (LDMs) (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022), which refine latent space representations through a diffusion process. LDMs iteratively denoise latent variables, progressively improving the quality of generated samples. While these models can produce realistic-looking data, they often generate new samples of one single modality at a time. Thus, it is difficult for them to generate multiple modalities using one generative model as they lack the physical consistency to the governing PDEs that describe the relationship between these modalities. Diffusion models have been applied to FWI by Wang et al. (Wang et al., 2023a), who used them to generate prior distributions for plausible velocity models as a regularization term. Their method still treats seismic data and velocity maps separately, limiting its ability to generate physically consistent seismic-velocity pairs.
5 Conclusion
In this work, we revisit FWI from a generative perspective and propose WaveDiffusion, a diffusion-based framework that refines the latent space constructed by encoder-decoder models. We demonstrate that standard image-to-image translation FWI models do not ensure physical consistency in their learned latent representations. To address this, we employ a joint diffusion model in the shared latent space that learns to score deviations from the governing PDE, progressively transforming arbitrary latent points into physically consistent solutions, ensuring that generated seismic-velocity pairs naturally satisfy the governing PDE. Additionally, our experiments show that the diffusion model can improve the encoder-decoder performance of FWI tasks, and the generated data can serve as the augmentation of the training set for data-driven FWI models, particularly in scenarios with limited real data availability. Our findings provide a new perspective on bridging generative modeling with physics-based problem-solving, paving the way for diffusion models to enhance scientific discovery and computational physics applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Network Details and Training Hyperparameters
In this appendix section, we provide details on the network architectures and training hyperparameters used for the encoder-decoder FWI model and joint diffusion models in our experiments.
A.1.1 encoder-decoder Model
The encoder and decoder use convolutional encoder-decoder branches that are constructed by ResNet blocks (He et al., 2016) for seismic and velocity data. The channel multipliers of the ResNet blocks are set to [1, 2, 2, 4, 4] for velocity maps and [1, 2, 2, 4, 4, 4, 4, 8, 8] for seismic data. The resolution for velocity maps is 64, while for seismic data, it is [1024, 64]. The size of the latent feature map $z$ is [16, 16], and the number of residual blocks is set to 3. The model was trained with a base learning rate of $4.5\times 10^{-4}$. It uses an embedding dimension of 32 and an embedding codebook size of 8192. It employs a perceptual loss combined with a discriminator. The discriminator starts training at step 50001 with a discriminator weight of 0.5 and a perceptual weight of 0.5.
A.1.2 Joint Diffusion Model
The Joint Diffusion model is based on the LatentDiffusion architecture. The backbone network in the Joint Diffusion model is a UNet-based architecture. The UNet takes 32 input and output channels, and the model channels are set to 128. The attention resolutions are [1, 2, 4, 4], corresponding to spatial resolutions of 32, 16, 8, and 4. The model uses 2 residual blocks and channel multipliers of [1, 2, 2, 4, 4]. It also employs 8 attention heads with scale-shift normalization enabled and residual blocks that support upsampling and downsampling. The model is trained with a base learning rate of $5.0\times 10^{-5}$ and uses 1000 diffusion timesteps. The loss function applied is $L_{1}$. The diffusion process is configured with a linear noise schedule, starting from 0.0015 and ending at 0.0155.
A LambdaLinearScheduler is used to control the learning rate, with 10000 warmup steps. The initial learning rate is set to $1.0\times 10^{-6}$, which increases to a maximum of $1.0$ over the course of training.
Figure 9: Generated Examples from the joint diffusion model with one VQ.
A.1.3 Training Hyperparameters
Both the encoder-decoder model and joint diffusion models were trained using the Adam optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The models were trained with a batch size of 16 for 500 epochs. The learning rate follows an exponential decay schedule with a decay rate of 0.98. Gradient clipping was applied with a threshold of 1.0. Early stopping was implemented when the validation loss plateaued for 10 consecutive epochs.
We trained our models on 128 NVIDIA GH200 GPUs. Training required approximately 8000 GPU hours for the first-stage encoder-decoder model and 12000 GPU hours for the joint diffusion model.
Seismic data and velocity models were resized from [5,70,1000]/[1,70,70] to [3,64,1024]/[1,64,64] (channel, height, depth) for consistency with our architecture. Log transform is performed for seismic data. Both were normalized to [-1,1] to ensure compatibility and stability.
A.2 Joint Generation Examples
We illustrate generated samples from the model with one VQ in Figure 9. In the top row, we present samples structurally similar to FlatVel-B (FVB), where the seismic inputs exhibit perfect symmetry along a central vertical plane. The corresponding generated velocity maps preserve this symmetry, demonstrating the model’s ability to respect the geometric constraints of the input data.
The model also generates samples that fuse structures from multiple datasets, as observed in the second row of Figure 9. These examples showcase that the model effectively handles a wide range of input patterns while maintaining physical and structural coherence.
A.3 Statistical Comparison of Our Models and BigFWI-B
We present the detailed statistical comparison between the BigFWI-B model and our first-stage encoder-decoder model and the corresponding diffusion models on the inversion tasks for the velocity maps across all the datasets of OpenFWI.
A.4 InversionNet Results Visulization
In Figure 11, we illustrate more visualizations of InversinNet Results trained on different training data settings.
A.5 Separate vs. Joint Diffusion
We compare the joint diffusion model with separate diffusion models, where seismic data and velocity maps are generated independently. In the separate models, the latent space constructed by the single-branch encoder-decoder lacks a shared representation. Both approaches are trained on the CVB subset, with results summarized in Table 4 and Figure 10.
The joint diffusion model consistently outperforms the separate models in FID scores. For the CVB dataset, joint diffusion achieves FID scores of 30.66 (seismic) and 186.86 (velocity), compared to 131.48 and 411.40, respectively, for the separate models. This highlights the superior quality of the joint diffusion outputs.
Table 4: FID score comparison between separate and joint generations. Evaluations on seismic data $s$ and velocity maps $v$ on the CVB dataset.
Modality LDM Setup FID
Velocity Joint 186.86
Separate 411.40
Seismic Joint 30.66
Separate 131.48
Beyond visual quality, the joint model enforces physical consistency with the wave equation, which the separate models fail to achieve. As shown in Figure 10, separate models exhibit significant deviations from the governing PDE, while the joint diffusion model generates seismic-velocity pairs that are both visually realistic and physically valid. This underscores the effectiveness of the WaveDiffusion framework in maintaining fidelity and physical correctness.
Figure 10: Visualization of separate vs. joint diffusion. Row 1: Separate models; Row 2: Joint model. Column 4 shows deviations from the governing PDE.
Table 5: Performance Comparison of Stage 1 Encoder-Decoders and Diffusion Models on OpenFWI Datasets. Metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Structural Similarity Index (SSIM). Lower values for MAE and RMSE indicate better performance, while higher SSIM values indicate better structural similarity.
Dataset Metric Model
BigFWI-B 2VQ 1VQ 2VQ LDM 1VQ LDM
FlatVel_A MAE 0.0055 0.0097 0.0197 0.0087 0.0198
RMSE 0.0130 0.0133 0.0254 0.0113 0.0222
SSIM 0.9943 0.9974 0.9943 0.9989 0.9958
FlatVel_B MAE 0.0233 0.0203 0.0396 0.0195 0.0402
RMSE 0.0696 0.0273 0.0619 0.0285 0.0560
SSIM 0.9658 0.9952 0.9746 0.9936 0.9776
CurveVel_A MAE 0.0343 0.0281 0.0383 0.0251 0.0315
RMSE 0.0798 0.0652 0.0603 0.0604 0.0585
SSIM 0.9027 0.9282 0.9401 0.9384 0.9451
CurveVel_B MAE 0.0933 0.0658 0.0794 0.0640 0.0733
RMSE 0.2154 0.1610 0.1523 0.1598 0.1519
SSIM 0.7808 0.8541 0.8590 0.8556 0.8633
FlatFault_A MAE 0.0106 0.0186 0.0197 0.0148 0.0186
RMSE 0.0286 0.0231 0.0279 0.0237 0.0262
SSIM 0.9871 0.9893 0.9871 0.9901 0.9893
Style_A MAE 0.0514 0.0553 0.0575 0.0462 0.0495
RMSE 0.0868 0.0730 0.0767 0.0694 0.0738
SSIM 0.9125 0.9334 0.9313 0.9384 0.9359
Style_B MAE 0.0553 0.0626 0.0637 0.0637 0.0630
RMSE 0.0876 0.0969 0.0947 0.0960 0.0944
SSIM 0.7567 0.7289 0.7368 0.7320 0.7381
Figure 11: InversionNet performance visualization. The predictions on the FVB, CVB, FFB, CFA, and SA subsets.